use node analysis to find I1, I2, I3 and I4

Use Node Analysis To Find I., I2 I3, IA IA V 6A LI. 200 172 & 1or 302 120A &GR
Answer #1

14 şior {300 D 200 şar } zor Applying kel at hode o. { leaving current zo -6+ x + = + = -20 + Ž=0 -360 +3V+5 +26 - 1200 + lor